OpenStudy (anonymous):

Why cannot we divide by 0

OpenStudy (anonymous):

Well, maybe textbooks need to stop putting 0 under numbers and confusing students.

OpenStudy (kinggeorge):

Basically, if you divide by 0, you can get a bunch of contradictions like 1=2. For example, if we say \(1/0=\infty\) and \(2/0=\infty\), then \(1=0\cdot\infty=2\) so \(1=2\).

OpenStudy (anonymous):

I thought 0 had no effect.?

OpenStudy (kinggeorge):

On a deeper level, if you divide by 0, it's basically equivalent to saying that there exists some number \(x\) such that \(0x=1\). Since this is obviously false for all real numbers \(x\), we can't divide by 0.

OpenStudy (zzr0ck3r):

think about these examples 1/.001 = big number 1/.0000001 = biger number 1/.000000000000000001 = crazy big number 1/.000000000000000000000000000000000000000000000000000000000001 = omg this number is huge 1/0 = the biggest number ever

OpenStudy (zzr0ck3r):

now, is there a biggest number ever?

OpenStudy (anonymous):

Oh my.

OpenStudy (nincompoop):

put it into words: divide something by nothing… you don't get anywhere, and that's why you wind up with undefined or complex infinity.

hero (hero):

There's a specific process that takes place when you divide. When you divide, what you're really doing is asking yourself, how many times can you subtract the denominator from the numerator in order to get to zero. For example, 16/4 = 4 because You can subtract 4 from 16 four times in order to get back to zero. 16 - 4 = 12 12 - 4 = 8 8 - 4 = 4 4 - 4 = 0 But if you have something like 16/0, you'll never get back to zero...you're sort of stuck with subtracting zero from 16 forever or unto infinity. The goal of division is to get back to zero. Dividing by zero doesn't accomplish that.

OpenStudy (tkhunny):

Division by zero does not create a number. It is certainly not a big number. In the Rational World 6/2 = 3 It takes 3 2's to get you to 6 15/3 = 5 It takes 5 3's to get you to 15 80/10 = 8 It takes 8 10's to get you to 80 5/0 = NaN This is no number of 0's that will get you to 5.